# ESIGMAHM: An Eccentric, Spinning inspiral-merger-ringdown waveform model with Higher Modes for the detection and characterization of binary black holes

### v1

@@ III Waveform Model Construction
The framework ESIGMA is built with two primary components. The first one deals with the inspiral evolution that combines input from post-Newtonian theory [146] and from the self-force approach [147, 148] describing the dynamics of spinning, eccentric compact binary systems. While the second is the MR segment that has been described by a NR based surrogate model. The following subsections elaborate on these individual pieces.
@@ III.1 Inspiral Model
The inspiral segment of ESIGMA framework is modeled using inputs from the PN theory and self-force program. These corrections are incorporated in the framework in two forms. First is in the evolution of orbital elements such as eccentricity $(e)$, PN parameter $(x)$, mean anomaly $(l)$, and orbital phase $(\phi)$ that describe the dynamical evolution of the system. The evolution equations for these orbital variables, including spins and eccentricity, were computed up to 3PN order in Ref. [100, 107]. The generalized quasi-Keplerian (QK) representation for non-spinning [149, 150, 141] and spinning systems [100, 107] are used to describe the inspiral evolution. Besides, 6PN corrections from the gravitational self-force approach to the binding energy of compact binaries and energy flux of quasi-circular binaries are also included. When modeling the inspiral we work with the adiabatic approximation in which radiation reaction time-scale is much longer compared to orbital and precessional time scales associated with eccentric orbits [151]. Under this assumption, we can use an orbit-averaged description of the variables, describing the conservative and radiative dynamics of the system.
@@
where the angles $\left(\Theta,\Phi\right)$ specify the binary’s orientation with respect to the observer at a reference. Both instantaneous (the part of GW radiation that depends on the state of the source at a given retarded time) and hereditary effects (the part of GW radiation which depends on the dynamical history of the source) [^3] Hereditary contributions include tails, tails-of-tails, tail-of-memory and the non-linear memory contributions in the gravitational waveform [146]. are included in the inspiral version of our model. The instantaneous terms in waveform modes include non-spinning and spinning corrections for general orbits, while the hereditary part only includes corrections in the quasi-circular orbit.
@@ III.2 Merger and Ringdown
Figure 1: As expected, post-inspiral circularisation of gravitational waveforms for nearly all spinning eccentric NR simulations listed in Table 3 is observed (grey curves); allowing for quasi-circular MR attachment at a frequency $f_{\rm{trans}}$. For comparison, a quasi-circular simulation is also plotted (brown curve) and since Kerr ISCO frequency is always larger than the Schwarzschild ISCO frequency for positive-aligned spins, solid circles (brown) are shown only for anti-aligned cases.
While post-Newtonian approximation accurately describes binary’s early inspiral dynamics, it breaks down towards the late inspiral stage as the system evolves into highly relativistic, non-perturbative regime. This later stage is most accurately described by solutions to Einstein’s equations within NR. Significant efforts have been made to develop NR waveforms including the effects of spins and eccentricity with equal-mass to high mass-ratio cases [175, 176, 177]. As in the ENIGMA framework, we assume that most of the orbital eccentricity has decayed by the end of the inspiral, and post-inspiral is indistinguishable from coalescences on quasi-circular orbits. We use NRSur7dq4 [91], a quasi-circular spinning model for merger and ringdown.
To construct the full IMR model, we extract individual spherical harmonic modes of the waveforms and combine the inspiral portion with the MR by creating an interpolant that stitches the amplitude and frequency of the modes in a smooth manner. The interpolated amplitude and frequency are then compiled to create the IMR modes, which are then summed over to generate the GW strain of Eq. (7).
This algorithm has three key steps: matching the transition time-window between inspiral and MR, aligning their phase in time domain, and applying the stitch. We start with specifying an attachment frequency $f_{\rm{trans}}$, and locate the times $t_{\rm{trans}}^{\rm{Insp}}$ and $t_{\rm{trans}}^{\rm{MR}}$ in the inspiral and MR waveforms where their orbital (or half of the dominant-mode) frequency crosses $f_{\rm{trans}}$. For the eccentric inspiral, we consider the orbit-averaged (and not the instantaneous) orbital frequency. These times are identified with each other, and denote the end of the transition time-window. We then compute the integral of the orbit-averaged frequency of the inspiral model backwards in time starting from $t_{\rm{trans}}^{\rm{Insp}}$, and locate the time where the integral crosses a specified value $\delta_{\rm{trans}}$. This is the time $t_{\delta}$ with:
$\phi(t_{\delta}):=\phi(t_{\rm{trans}}^{\rm{Insp}})-\delta_{\rm{trans}},$ (8)
@@
The start of the transition windows then correspond to the times $t_{\rm{trans,0}}^{\rm{Insp}}$ $=$ $t_{\rm{trans}}^{\rm{Insp}}$ $-t_{\delta}$ and $t_{\rm{trans,0}}^{\rm{MR}}$ $=$ $t_{\rm{trans}}^{\rm{MR}}$ $-t_{\delta}$ in the inspiral and MR waveforms. We find that setting $f_{\rm{trans}}$ to the minimum of Schwarzschild innermost stable circular orbit (ISCO) frequency and spinning Kerr ISCO frequency [178] works well in the quasi-circular limit. Similarly, we set $\delta_{\rm{trans}}=\pi/2$ radians to set the transition window to be a quarter orbital cycle in length.
Next, we align the inspiral and MR portions by minimizing their mismatch within this transition time-window. We add a constant phase offset $\phi_{a}$ to the phase of the MR modes $h_{\rm MR}^{\ell m}$ and align them with inspiral by minimizing the following mismatch measure $\Delta(\phi_{a})$ over $\phi_{a}$:
$\Delta(\phi_{a})=\frac{||h^{\ell m}_{\rm{Insp}}-h^{\ell m}_{\rm{MR}}e^{\mathrm{i}m\phi_{a}}||^{2}}{||h^{\ell m}_{\rm{MR}}||^{2}}$ (10)
We use the dominant mode with $\ell=m=2$ to find $\phi_{a}$ to perform this phase alignment, and phase-shift other modes $h^{\ell^{\prime}m^{\prime}}$ by $(m^{\prime}/2)\,\phi_{a}$ radians. This is motivated by identifying $\phi_{a}$ as twice the offset needed to align the orbital frequency evolutions between inspiral and MR. We compared the accuracy of the phase alignment between the following two approaches: minimizing the mismatch in Eq. 10 between just the dominant mode of inspiral and MR waveforms, and between the full inspiral and MR polarizations constructed by summing over several sub-dominant modes in addition to the dominant ones. The two approaches showed no significant difference in phase alignment, justifying our choice of using $h^{22}$ mode only.
@@
$A_{\rm{IMR}}^{\ell m}(t)=\begin{cases}A_{\rm{Insp}}^{\ell m}(t)&t<t_{i}\\ A_{\rm{Interp}}^{\ell m}(t)&t_{i}\leq t\leq t_{f}\\ A_{\rm{MR}}^{\ell m}&t_{f}<t\par\end{cases},$
$\phi_{\rm{IMR}}^{\ell m}(t)=\begin{cases}\phi_{\rm{Insp}}^{\ell m}(t)&t<t_{i}\\ \phi_{\rm{Interp}}^{\ell m}(t)&t_{i}\leq t\leq t_{f}\\ \phi_{\rm{MR}}^{\ell m}+\delta_{f}^{\ell m}&t_{f}<t\par\end{cases}.$
@@ III.3 Changes from previous versions
Figure 3: Agreement between InspiralESIGMA with existing EOB and Phenom waveforms are shown in quasi-circular limit. Match between quasi-circular dominant mode InspiralESIGMA waveforms with SEOBNRv4, SEOBNRv5_ROM and IMRPhenomXAS are shown for fixed total mass $40\,M_{\odot}$ and dimensionless spins $(0.5,0.5)$, respectively. The match computation is performed with a high-frequency cutoff of a minimum of ISCO frequencies between Schwarzschild and spinning cases [178] and using aLIGO zero-detuned high-power PSD assuming initial gravitational wave frequency of $10$Hz. The horizontal dashed black line denote the 99% match between the two waveforms.
@@
Figure 5: Comparison of strain data $(h_{+})$ of IMRESIGMA with 32 out of 42 (see Table 2) quasi-circular NR simulations taken from SXS catalog to validate IMRESIGMA in the zero-eccentricity limit. The comparison is done by including the dominant $\ell=|m|=2$ modes and using a total mass of $40M_{\odot}$. The inset shows a zoomed in view of the late inspiral and MR portion. The value of reference mass-ratio $(q)$, and the $z$-component of the reference dimensionless spins $(\chi_{1_{z}},\chi_{2_{z}})$ of each simulations are given in each figure.
Figure 6: Comparison of strain data $(h_{+})$ of IMRESIGMA with spinning, eccentric NR simulations taken from SXS catalog and listed in Table 3. The comparison is done by including $(\ell,|m|)=(2,1),(3,3),(3,2),(4,4),(4,3)$ in addition to $(2,2)$ mode and using a total mass of $40M_{\odot}$. The inset shows a zoomed in view of the late inspiral and MR portion. The mass ratio $(q)$, $z$-components of dimensionless spins $(\chi_{1_{z}},\chi_{2_{z}})$, and reference eccentricity $(e_{\rm{ref}})$ are given on the figure corresponding to each simulation. Simulations for which the eccentricity is not well measured, we quote the optimized eccentricity for those cases (denoted as $e^{\ast}_{\rm{ref}})$.
@@ III.4 Validation of model in the quasi-circular limit
In Fig. 3, we show matches between InspiralESIGMA and SEOBNRv4 [109], IMRPhenomXAS [182], and SEOBNRv5_ROM [112] in spinning quasi-circular limit for binaries with a fixed total mass $40M_{\odot}$ having dimensionless spins $(0.5,0.5)$ using the dominant $\ell=|m|=2$ modes. We fix the lower cutoff frequency at $10$Hz for all cases and take the upper cutoff frequency of the match calculation to be the minimum frequency between the ISCO frequencies for Schwarzschild (i.e., non-spinning) and spinning cases [178]. [^5] This choice corresponds to the inspiral to MR transition frequency ($f_{\rm{trans}}$) for IMRESIGMA (see Sec. III.2). From the figure, we see that InspiralESIGMAHM successfully reproduces the true dynamics of spinning systems in quasi-circular limit with match values $\mathcal{O}\gtrsim 99\%$ with both EOB and Phenom waveforms for mass-ratios up to $q=8$.
We next validate the full IMR model using quasi-circular NR waveforms. We choose a set of simulations with mass ratios between 1 and 5, and component spins (anti-) aligned with the orbital angular momentum from the SXS waveforms catalog [137]. In Fig. 5 we show IMRESIGMA strain data $(h_{+})$ versus quasi-circular NR simulations for $32$ out of $42$ non-spinning and aligned-spin cases (see Table 2 of Appendix D). As before, we fix the source total mass to $40M_{\odot}$ and use the dominant $\ell=|m|=2$ modes. The individual panels show the mass-ratio and component spins for each of the simulations. We notice that for both non-spinning and spinning systems, IMRESIGMA reproduces NR simulations very closely. Quantifying this further, we next compute matches between our model and the same set of NR simulations. We do so for a range of total masses for each simulation, and show the resulting mismatches in Fig. 4. Dashed horizontal lines mark $1\%,3\%$, and $10\%$ mismatches. We note that the agreement remains better than $3\%$ for most simulations, and better than $1\%$ for a majority of them. The largest mismatches $(3-7\%)$ are shown by a cluster of curves corresponding to binary configurations featuring anti-aligned component spins and unequal mass ratios. Other than for this corner of the binary black hole parameter space, we find IMRESIGMA to be largely faithful to NR simulations in the quasi-circular limit.
@@ IV Validation of model with eccentric Numerical Relativity Simulations
Next, Fig. 8 (left panel) shows that mismatches with NR simulations that have one or both the spins anti-aligned with the orbital angular momentum vector (SXS:BBH:0089, SXS:BBH:0105, SXS:BBH:0108, SXS:BBH:1136, SXS:BBH:1169) can be $>3\%$ in the mass range $20-120M_{\odot}$. Further, the mismatches are higher (can be as high as $10\%$) for the large mass ratio cases (SXS:BBH:0105, SXS:BBH:0108, SXS:BBH:1169) compared to others (SXS:BBH:0089, SXS:BBH:1136), as indicated by the right panel of the same figure which shows the highest mismatches to be for $q\gtrsim 3$ and $\chi_{\rm{eff}}<-0.3$. For simulations with larger initial eccentricity (such as SXS:BBH:1136), this is likely due to weakening of the assumption discussed in Sec. III that binary has circularized beyond the inspiral-merger transition frequency, which is smaller (and earlier) in case of anti-aligned spins as opposed to aligned spins. As a consequence, the attachment of eccentric inspiral with quasi-circular MR is happening at a point where the binary hasn’t sufficiently circularized yet. One remedy to this could be to use an eccentric MR model instead of a quasi-circular one. This would accurately represent the dynamics of the eccentric signal throughout the merger and ringdown. As a consequence, the match will improve for these NR simulations. For simulations with smaller initial eccentricity (such as SXS:BBH:0105) this could indicate the need to include higher-order spin information in ESIGMA. A remedy for this would be to include higher order spin-eccentricity dependent terms in the radiative sector and transition from inspiral to merger prescription closer to merger. We defer a detailed investigation to future work, and recommend the use of IMRESIGMAHM to model binaries of moderately spinning black holes merging on moderately eccentric orbits.

### v2

@@ III Waveform Model Construction
The framework ESIGMA is built with two primary components. The first one deals with the inspiral evolution that combines input from post-Newtonian theory [179] and from the self-force approach [180, 181] describing the dynamics of spinning, eccentric compact binary systems. While the second is the PMR segment that has been described by an NR based surrogate model. The following subsections elaborate on these individual pieces.
@@ III.1 Inspiral Model
The inspiral segment of ESIGMA framework is modeled using inputs from the PN theory and self-force program. These corrections are incorporated in the framework in two forms. First is in the evolution of orbital elements such as eccentricity $(e)$, PN parameter $(x)$, mean anomaly $(l)$, and orbital phase $(\phi)$ that describe the dynamical evolution of the system. The evolution equations for these orbital variables, including spins and eccentricity, were computed up to 3PN order in Ref. [130, 138]. The generalized quasi-Keplerian (QK) representation for non-spinning [182, 183, 174] and spinning systems [130, 138] are used to describe the inspiral evolution. Besides, 4PN corrections from the gravitational self-force approach to the binding energy of compact binaries and energy flux of quasi-circular binaries are also included. When modeling the inspiral we work with the adiabatic approximation in which radiation reaction time-scale is much longer compared to orbital and precessional time scales associated with eccentric orbits [184]. Under this assumption, we can use an orbit-averaged description of the variables, describing the conservative and radiative dynamics of the system.
@@
where the angles $\left(\Theta,\Phi\right)$ specify the binary’s orientation with respect to the observer at a reference. Both instantaneous (the part of GW radiation that depends on the state of the source at a given retarded time) and hereditary effects (the part of GW radiation which depends on the dynamical history of the source) included in the inspiral version of our model. The instantaneous terms in waveform modes include non-spinning and spinning corrections for general orbits, while the hereditary part only includes corrections in the quasi-circular orbit. [^3] Hereditary contributions include tails, tails-of-tails, tail-of-memory and the non-linear memory contributions in the gravitational waveform [179]. Note that $(\ell,0)$ modes are not included in the inspiral part of the model. Consequently, the non-linear Christodoulou memory [195] is not taken into account. The tail-of-memory and non-linear memory (except Christodoulou memory) contributions are present in modes other than $(\ell,0)$.
@@ III.2 Merger and Ringdown
Figure 1: As expected, post-inspiral circularisation of gravitational waveforms for nearly all spinning eccentric NR simulations listed in Table 3 is observed (grey curves); allowing for quasi-circular PMR attachment at a frequency $f_{\rm{trans}}$. For comparison, a quasi-circular simulation is also plotted (brown curve) and since Kerr ISCO frequency is always larger than the Schwarzschild ISCO frequency for positive-aligned spins, solid circles (brown) are shown only for anti-aligned cases.
While post-Newtonian approximation accurately describes binary’s early inspiral dynamics, it breaks down toward the late inspiral stage as the system evolves into highly relativistic, non-perturbative regime. This later stage is most accurately described by solutions to Einstein’s equations within NR. Significant efforts have been made to develop NR waveforms including the effects of spins and eccentricity with equal-mass to high mass-ratio cases [209, 210, 211]. As in the ENIGMA framework, we assume that most of the orbital eccentricity has decayed by the end of the inspiral, and post-inspiral is indistinguishable from coalescences on quasi-circular orbits. We use NRSur7dq4 [92], a quasi-circular spinning model for merger and ringdown.
To construct the full IMR model, we extract individual spherical harmonic modes of the waveforms and combine the inspiral portion with the PMR by creating an interpolant that stitches the amplitude and frequency of the modes in a smooth manner. The interpolated amplitude and frequency are then compiled to create the IMR modes, which are then summed over to generate the GW strain of Eq. (7).
This algorithm has three key steps: matching the transition time-window between inspiral and PMR, aligning their phase in time domain, and applying the stitch. We start with specifying an attachment frequency $f_{\rm{trans}}$, and locate the times $t_{\rm{trans}}^{\rm{Insp}}$ and $t_{\rm{trans}}^{\rm{PMR}}$ in the inspiral and PMR waveforms where their orbital (or half of the dominant-mode) frequency crosses $f_{\rm{trans}}$. For the eccentric inspiral, we consider the orbit-averaged (and not the instantaneous) orbital frequency. These times are identified with each other, and denote the end of the transition time-window. We then compute the integral of the orbit-averaged frequency of the inspiral model backward in time starting from $t_{\rm{trans}}^{\rm{Insp}}$, and locate the time where the integral crosses a specified value $\delta_{\rm{trans}}$. This is the time $t_{\delta}$ with:
$\phi(t_{\delta}):=\phi(t_{\rm{trans}}^{\rm{Insp}})-2\pi\delta_{\rm{trans}},$ (8)
@@
The start of the transition windows then correspond to the times $t_{\rm{trans,0}}^{\rm{Insp}}$ $=$ $t_{\rm{trans}}^{\rm{Insp}}$ $-t_{\delta}$ and $t_{\rm{trans,0}}^{\rm{PMR}}$ $=$ $t_{\rm{trans}}^{\rm{PMR}}$ $-t_{\delta}$ in the inspiral and PMR waveforms. We find that setting $f_{\rm{trans}}$ to the minimum of Schwarzschild innermost stable circular orbit (ISCO) frequency and spinning Kerr ISCO frequency [212] [^4] In order to obtain the Kerr ISCO frequency, we first obtain the dimensionless spin $a^{\rm{eff}}_{f}\left(\nu,S\right)$ of the Kerr remnant for the binary using Eq. (3.6) of Ref. [212], where $\nu$ is the dimensionless mass-ratio, and $S$ is the total spin of the binary. The ISCO frequency is calculated using the prescription in Eq. (2.23) (and appendix C) of Ref. [213]. works well in the quasi-circular limit  [^5] Note that these choices of transition frequency were arrived at by comparison with NR/EOB models in the quasicircular limit. For anti-aligned spin binaries our prescription for transition from the eccentric inspiral to quasicircular NR surrogate implies that it is done at a relatively earlier frequency than for aligned-spin configurations. This naturally results in ESIGMAHM limiting itself to relatively lower values of initial eccentricity for anti-aligned spin binaries. We plan to improve the inspiral model in the late-inspiral in a follow-up work, which will allow the transition to be done closer to merger.. Similarly, we set $\delta_{\rm{trans}}=\pi/2$ to set the transition window to be about 1.5 orbital cycles in length. The transition windows are shown as grey bands in all panels of Figs. 5 and 6.
Next, we align the inspiral and PMR portions by minimizing their mismatch within this transition time-window. We add a constant phase offset $\phi_{a}$ to the phase of the PMR modes $h_{\rm PMR}^{\ell m}$ and align them with inspiral by minimizing the following mismatch measure $\Delta(\phi_{a})$ over $\phi_{a}$:
$\Delta(\phi_{a})=\frac{||h^{\ell m}_{\rm{Insp}}-h^{\ell m}_{\rm{PMR}}e^{\mathrm{i}m\phi_{a}}||^{2}}{||h^{\ell m}_{\rm{PMR}}||^{2}}$ (10)
We use the dominant mode with $\ell=m=2$ to find $\phi_{a}$ to perform this phase alignment, and phase-shift other modes $h^{\ell^{\prime}m^{\prime}}$ by $(m^{\prime}/2)\,\phi_{a}$ radians. This is motivated by identifying $\phi_{a}$ as twice the offset needed to align the orbital frequency evolutions between inspiral and PMR. We compared the accuracy of the phase alignment between the following two approaches: minimizing the mismatch in Eq. (10) between just the dominant mode of inspiral and PMR waveforms, and between the full inspiral and PMR polarizations constructed by summing over several sub-dominant modes in addition to the dominant ones. The two approaches showed no significant difference in phase alignment, justifying our choice of using $h^{22}$ mode only.
@@
$A_{\rm{IMR}}^{\ell m}(t)=\begin{cases}A_{\rm{Insp}}^{\ell m}(t)&t<t_{i}\\ A_{\rm{Interp}}^{\ell m}(t)&t_{i}\leq t\leq t_{f}\\ A_{\rm{PMR}}^{\ell m}&t_{f}<t\par\end{cases},$
$\phi_{\rm{IMR}}^{\ell m}(t)=\begin{cases}\phi_{\rm{Insp}}^{\ell m}(t)&t<t_{i}\\ \phi_{\rm{Interp}}^{\ell m}(t)&t_{i}\leq t\leq t_{f}\\ \phi_{\rm{PMR}}^{\ell m}+\delta_{f}^{\ell m}&t_{f}<t\par\end{cases}.$
@@ III.3 Changes from previous versions
Figure 3: Agreement between InspiralESIGMA with existing EOB and Phenom waveforms are shown in quasi-circular limit. Match between quasi-circular dominant mode InspiralESIGMA waveforms with SEOBNRv4, SEOBNRv5_ROM and IMRPhenomXAS are shown for fixed total mass $40\,M_{\odot}$ and dimensionless spins $(0.5,0.5)$, respectively. The match computation is performed with a high-frequency cutoff of a minimum of ISCO frequencies between Schwarzschild and spinning cases [212, 213] and using aLIGO zero-detuned high-power PSD assuming initial gravitational wave frequency of $10$Hz. The horizontal dashed black line denote the 99% match between the two waveforms.
@@
Figure 5: Comparison of strain data $(h_{+})$ of IMRESIGMA with 32 out of 42 (see Table 2) quasi-circular NR simulations taken from SXS catalog to validate IMRESIGMA in the zero-eccentricity limit. The comparison is done by including the dominant $\ell=|m|=2$ modes and using a total mass of $40M_{\odot}$. The inset shows a zoomed in view of the late inspiral and PMR portion. The gray shaded region denotes the time interval over which the IMRESIGMA transitions from its inspiral prescription to the plunge-merger-ringdown prescription. The value of reference mass-ratio $(q)$, and the $z$-component of the reference dimensionless spins $(\chi_{1_{z}},\chi_{2_{z}})$ of each simulations are given in each figure.
Figure 6: Comparison of strain data $(h_{+})$ of IMRESIGMA with spinning, eccentric NR simulations taken from SXS catalog and listed in Table 3. The comparison is done by including $(\ell,|m|)=(2,1),(3,3),(3,2),(4,4),(4,3)$ in addition to $(2,2)$ mode and using a total mass of $40M_{\odot}$. The inset shows a zoomed in view of the late inspiral and PMR portion. The gray shaded region denotes the time interval over which the IMRESIGMA transitions from its inspiral prescription to the plunge-merger-ringdown prescription. The mass ratio $(q)$, $z$-components of dimensionless spins $(\chi_{1_{z}},\chi_{2_{z}})$, and reference eccentricity $(e_{\rm{ref}})$ are given on the figure corresponding to each simulation. Simulations for which the eccentricity is not well measured, we quote the optimized eccentricity for those cases (denoted as $e^{\ast}_{\rm{ref}})$.
@@ III.4 Validation of model in the quasi-circular limit
In Fig. 3, we show matches between InspiralESIGMA and SEOBNRv4 [100], IMRPhenomXAS [217], and SEOBNRv5_ROM [118] in spinning quasi-circular limit for binaries with a fixed total mass $40M_{\odot}$ having dimensionless spins $(0.5,0.5)$ using the dominant $\ell=|m|=2$ modes. We fix the lower cutoff frequency at $10$Hz for all cases and take the upper cutoff frequency of the match calculation to be the minimum frequency between the ISCO frequencies for Schwarzschild (i.e., non-spinning) and spinning cases [212]. [^7] This choice corresponds to the inspiral to PMR transition frequency ($f_{\rm{trans}}$) for IMRESIGMA (see Sec. III.2). From the figure, we see that InspiralESIGMAHM successfully reproduces the true dynamics of spinning systems in quasi-circular limit with match values $\mathcal{O}\gtrsim 99\%$ with both EOB and Phenom waveforms for mass-ratios up to $q=8$.
We next validate the full IMR model using quasi-circular NR waveforms. We choose a set of simulations with mass ratios between 1 and 5, and component spins (anti-) aligned with the orbital angular momentum from the SXS waveforms catalog [169]. In Fig. 5 we show IMRESIGMA strain data $(h_{+})$ versus quasi-circular NR simulations for $32$ out of $42$ non-spinning and aligned-spin cases (see Table 2 of Appendix D). As before, we fix the source total mass to $40M_{\odot}$ and use the dominant $\ell=|m|=2$ modes. The individual panels show the mass-ratio and component spins for each of the simulations. We notice that for both non-spinning and spinning systems, IMRESIGMA reproduces NR simulations very closely. Quantifying this further, we next compute matches between our model and the same set of NR simulations. We do so for a range of total masses for each simulation, and show the resulting mismatches in Fig. 4. Dashed horizontal lines mark $1\%,3\%$, and $10\%$ mismatches. We note that the agreement remains better than $3\%$ for most simulations, and better than $1\%$ for a majority of them. The largest mismatches $(3-7\%)$ are shown by a cluster of curves corresponding to binary configurations featuring anti-aligned component spins and unequal mass ratios. Other than for this corner of the binary black hole parameter space, we find IMRESIGMA to be largely faithful to NR simulations in the quasi-circular limit. It is important to note while comparing IMRESIGMA with quasi-circular NR simulations, we found that the best fit IMRESIGMA waveform (that has the highest match with the NR simulation as given in Table 2) has a very small $(\sim 10^{-3})$ non-zero value of eccentricity. This can likely be fixed by including higher-order post-adiabatic initial conditions which we plan to explore in a future version of IMRESIGMA.
@@ IV Validation of model with eccentric Numerical Relativity Simulations
Next, Fig. 8 (left panel) shows that mismatches with NR simulations that have one or both the spins anti-aligned with the orbital angular momentum vector (SXS:BBH:0089, SXS:BBH:0105, SXS:BBH:0108, SXS:BBH:1136, SXS:BBH:1169) can be $>3\%$ in the mass range $20-120M_{\odot}$. Further, the mismatches are higher (can be as high as $10\%$) for the large mass ratio cases (SXS:BBH:0105, SXS:BBH:0108, SXS:BBH:1169) compared to others (SXS:BBH:0089, SXS:BBH:1136), as indicated by the right panel of the same figure which shows the highest mismatches to be for $q\gtrsim 3$ and $\chi_{\rm{eff}}<-0.3$. For simulations with larger initial eccentricity (such as SXS:BBH:1136), this is likely due to weakening of the assumption discussed in Sec. III that binary has circularised beyond the inspiral-merger transition frequency, which is smaller (and earlier) in case of anti-aligned spins as opposed to aligned spins. As a consequence, the attachment of eccentric inspiral with quasi-circular PMR is happening at a point where the binary hasn’t sufficiently circularized yet. One remedy to this could be to use an eccentric PMR model instead of a quasi-circular one. This would accurately represent the dynamics of the eccentric signal throughout the merger and ringdown. As a consequence, the match will improve for these NR simulations. For simulations with smaller initial eccentricity (such as SXS:BBH:0105) this could indicate the need to include higher-order spin information in ESIGMA. A remedy for this would be to include higher order spin-eccentricity dependent terms in the radiative sector and transition from inspiral to merger prescription closer to merger. We defer a detailed investigation to future work, and recommend the use of IMRESIGMAHM to model binaries of moderately spinning black holes merging on moderately eccentric orbits.